OpenStudy (judson_rockets):

What is the surface area of this prism? A. 1656 m2 B. 6480 m2 C. 3240 m2 D. 3312 m2

OpenStudy (abdullahm):

First find the area of the two triangles. Then find the area of the three rectangles. And then add up all of the areas.

OpenStudy (abdullahm):

The two rectangles that are slanted have the same area. The rectangle on the bottom has a different area.

OpenStudy (judson_rockets):

Caint do it

OpenStudy (abdullahm):

Let's find the area of the two triangles. They both have the same area. Area of a triangle = 1/2 * base * height Can you tell me the base and height of the triangle?

OpenStudy (judson_rockets):

The height is 12 And the base is 30

OpenStudy (judson_rockets):

I think

OpenStudy (abdullahm):

No, look carefully. The height is 12. But what is the base?

OpenStudy (judson_rockets):

It is 18

OpenStudy (abdullahm):

Correct. Now we have two triangles. Area = 1/2 * 18 * 12 What is the area?

OpenStudy (judson_rockets):

Well 18*12 is 216 but how much is a Half

OpenStudy (abdullahm):

Multiplying by 1/2 is the same as dividing by 2. So what is 216/2 = ?

OpenStudy (judson_rockets):

108

OpenStudy (abdullahm):

Good. Now we have two triangles. That means the area of one triangle is 108, and the other is also 108. 108 + 108 = 2(108) = ??

OpenStudy (judson_rockets):

216

OpenStudy (abdullahm):

So that's the area of the triangles. Now, we need to find the area of the rectangles. There are two rectangles with the same area. Let's take care of those first. The length is 30 and the width is 15. What is the area?

OpenStudy (judson_rockets):

450

OpenStudy (abdullahm):

And we have two triangles with that area. What is 450 + 450 = ?

OpenStudy (judson_rockets):

900

OpenStudy (abdullahm):

Ok, so we have the area of two rectangles and two triangles. Area of 2 triangles = 216 Area of 2 rectangles = 900 Total area so far is 1116 Let's find the area of the last rectangle now! The length is 30, and the width is 18. What is the area?

OpenStudy (judson_rockets):

540

OpenStudy (abdullahm):

And what is 1116 + 540 = ?

OpenStudy (judson_rockets):

1656

OpenStudy (judson_rockets):

Thank you

OpenStudy (abdullahm):

It was my pleasure!